# Relative topological principality and the ideal intersection property for groupoid C*-algebras

### v1

@@ 6. Examples from integer arithmetic
It is straightforward to verify (6.2) using the composition rule, see e.g. [Laca-Schulz23], proving the right LCM property. Notice that the principal right ideal $(a,[k])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ does not depend on $[k]$. Since ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ is countable, second countability of $\mathscr{G}$ follows from Lemma 5.8 and its proof. The only thing left to show is that $\mathscr{G}$ is Hausdorff. Recall from [SpielbergSmallCats, Definition 3.2.] that a left cancellative small category $\mathcal{C}$ is finitely aligned if for every $a,b\in\mathcal{C}$ there exists a finite set $F\subseteq\mathcal{C}$ such $a\mathcal{C}\cap b\mathcal{C}=\bigcup_{c\in F}c\mathcal{C}$. Right LCM monoids are obviously finitely aligned as small categories, with $F=\{c\}$. By [LiGarsideCats, Corollary 4.2.], to prove $\mathscr{G}$ is Hausdorff it suffices to show that for all $(a,[k]),(b,[l])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ the set
@@
A similar result to Proposition 6.6 can be proved about the monoid $\mathbb{N}^{\times}\ltimes\mathbb{Z}$, which is group embeddable, and $\mathbb{N}^{\times}\ltimes\mathbb{Q}/\mathbb{Z}$, which is not.
In contrast to the above applications we now consider the submonoid $\mathbb{Z}\rtimes\mathbb{N}^{\times}$ of the group $\mathbb{Q}\rtimes\mathbb{Q}_{+}^{*}$, where the operation is given by $(k,a)(l,b)=(k+al,ab)$ and the set of invertible elements is (×)*=$(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}=\mathbb{Z}\rtimes\{1\}$. This monoid is the opposite of $\mathbb{N}^{\times}\ltimes\mathbb{Z}$ and corresponds to the additive boundary quotient in [LacaRaeburn]. It has a finer principal ideal structure, see [LacaRaeburn, Section 5] and the isotropy at principal characters is consequently smaller. As a result, we see that ideal detection is more efficient using the interior of the isotropy, which in this case is simply the unit space.
@@
The action of $(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}$ on $\Omega$ is topologically free by [ToeLaNe, Lemma 1.6], and by [LS, Theorem 5.9] so is the partial action of $\mathbb{Q}_{+}^{*}\rtimes\mathbb{Q}$. Thus the interior of the isotropy is the unit space, which is properly contained in the action subgroupoid (×)*$(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$. For example, the element $((1,1),\chi_{(0,2)})\in(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$ is not in $\operatorname{Iso}(\mathscr{G})$ because $(1,1)\cdot\nolinebreak\chi_{(0,2)}=\chi_{(1,2)}\neq\chi_{(0,2)}$. ∎

### v3

@@ 6. Examples from integer arithmetic
It is straightforward to verify (6.2) using the composition rule, see e.g. [Laca-Schulz23], proving the right LCM property. Notice that the principal right ideal $(a,[k])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ does not depend on $[k]$.
Since ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ is countable, second countability of $\mathscr{G}$ follows from Lemma 5.9 and its proof. The only thing left to show is that $\mathscr{G}$ is Hausdorff. Recall from [SpielbergSmallCats, Definition 3.2.] that a left cancellative small category $\mathcal{C}$ is finitely aligned if for every $a,b\in\mathcal{C}$ there exists a finite set $F\subseteq\mathcal{C}$ such $a\mathcal{C}\cap b\mathcal{C}=\bigcup_{c\in F}c\mathcal{C}$. Right LCM monoids are obviously finitely aligned as small categories, with $F=\{c\}$. By [LiGarsideCats, Corollary 4.2.], to prove $\mathscr{G}$ is Hausdorff it suffices to show that for all $(a,[k]),(b,[l])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ the set
@@
From the proof of Proposition 6.6 we see that (×$(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega$ is an open subgroupoid of the interior of the isotropy. Moreover, at the canonical dense subset of characters $\chi_{(a,[m])}$ the isotropy of this subgroupoid coincides with that of $\mathscr{G}$. Thus, our Corollary 6.3 already follows from Starling’s improvement [Starling, Theorem 2.1(ii)] of the criteria from [brown2015cartan].
A similar result to Proposition 6.6 can be proved about the monoid $\mathbb{N}^{\times}\ltimes\mathbb{Z}$, which is group embeddable, and about $\mathbb{N}^{\times}\ltimes\mathbb{Q}/\mathbb{Z}$, which is not.
In contrast to the above applications we now consider the submonoid $\mathbb{Z}\rtimes\mathbb{N}^{\times}$ of the group $\mathbb{Q}\rtimes\mathbb{Q}_{+}^{*}$, where the operation is given by $(k,a)(l,b)=(k+al,ab)$ and the set of invertible elements is (×)=$(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}=\mathbb{Z}\rtimes\{1\}$. This monoid is the opposite of $\mathbb{N}^{\times}\ltimes\mathbb{Z}$ and corresponds to the additive boundary quotient in [LacaRaeburn]. It has a finer principal ideal structure, see [LacaRaeburn, Section 5] and the isotropy at principal characters is consequently smaller. As a result, we see that ideal detection is more efficient using the interior of the isotropy, which in this case is simply the unit space.
@@
The action of $(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}$ on $\Omega$ is topologically free by [ToeLaNe, Lemma 1.6], and by [LS, Theorem 5.9] so is the partial action of $\mathbb{Q}_{+}^{*}\rtimes\mathbb{Q}$. Thus the interior of the isotropy is the unit space, which is properly contained in the action subgroupoid (×$(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$. For example, the element $((1,1),\chi_{(0,2)})\in(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$ is not in $\operatorname{Iso}(\mathscr{G})$ because $(1,1)\cdot\nolinebreak\chi_{(0,2)}=\chi_{(1,2)}\neq\chi_{(0,2)}$. ∎
Disclosure and data availability statement
The authors have no relevant financial or non-financial interests to disclose. There is no data associated to this paper.